OpenStudy (naudiapie):

ennifer and Taylor are running together. Taylor thinks Jennifer runs a little slow, so she gives Jennifer a 1 km advantage. This means Taylor is at the starting point and Jennifer is 1 km in front. Taylor runs 0.25 km per minute and Jennifer runs 0.15 km. How long will it take before Taylor catches up to Jennifer?

OpenStudy (naudiapie):

helppp pls

OpenStudy (mathstudent55):

What is the difference in their speeds?

OpenStudy (naudiapie):

0.10?

OpenStudy (mathstudent55):

Correct. The difference in speed is 0.10 km/minute. How many minutes will it take for Taylor to run 1 km more than Jennifer being that Taylor runs 0.10 km/minute faster?

OpenStudy (naudiapie):

not sure

OpenStudy (naudiapie):

is there formula??

OpenStudy (mathstudent55):

Time (min) 0 1 2 3 4 5 6 7 8 9 10 Taylor (km) 1 1.15 1.30 1.45 1.60 1.75 1.90 2.05 2.20 2.35 2.5 Jennifer (km) 0 0.25 0.50 0.75 1.00 1.25 1.50 1.75 2.00 2.25 2.5

OpenStudy (mathstudent55):

We can write an equation. Also, we can reason it out. Taylor runs 0.10 km/minute faster than Jennifer. In 10 minutes, Taylor will have run 1 km more than Jennifer. Since Jennifer starts 1 km ahead, after 10 minutes, Taylor catches her.

OpenStudy (naudiapie):

so the answer is 10 min?

OpenStudy (mathstudent55):

Yes.

OpenStudy (mathstudent55):

I'll show it to you with an equation below.

OpenStudy (naudiapie):

did you get it by dividing 1 by 0.10

OpenStudy (mathstudent55):

distance time speed Jennifer x - 1 t 0.15 Taylor x t 0.25 Taylor runs a total of x distance. Since Jennifer has a 1 km head start, she runs 1 km less than Taylor, so Jennifer runs x - 1 distance. They both take t time until they are side by side. The speeds are 0.15 for Jennifer and 0.25 for Taylor. speed = distance/time to solve for time, multiply both sides by time and divide both sides by speed time = distance/speed Now we write an equation for each person using the equation just above and the info from the table Jennifer: t = (x - 1)/0.15 Taylor: t = x/0.25 Since t in both equations is equal, we can equate the equations: (x - 1)/0.15 = x/0.25 \(\dfrac{x - 1}{0.15} = \dfrac{x}{0.25} \) Cross multiply: 0.25(x - 1) = 0.15x 0.25x - 0.25 = 0.15x 0.1x - 0.25 = 0 0.1x = 0.25 x = 2.5 They meet at distance 2.5 km from the start. Now we can find the time. t = x/0.25 t = 2.5/0.25 t = 10 They meet at 10 minutes form the start.

OpenStudy (mathstudent55):

You are correct. Taylor has to make up 1 km of distance. Since the difference in speeds is 0.10 km/min, when you divide distance by speed you get time, so (1 km)/(0.1 km/min) = 10 minutes

OpenStudy (naudiapie):

So you can do it with 2 equations>

OpenStudy (mathstudent55):

We solved this problem three different ways. We did it by seeing where they both are at the end of each minute until their distances are the same. We also did it by reasoning and looking at the difference in speeds. Then we did it with equations.

OpenStudy (mathstudent55):

Yes. Use the two equations as a system of equations.

OpenStudy (naudiapie):

i find it easier to simply divide 1 by 0.1

OpenStudy (naudiapie):

thank you for the help

OpenStudy (mathstudent55):

You can set up two equations like we did above, or you can say: speed = distance/time Jennifer: 0.15 = (x-1)/t Taylor: 0.25 = x/t Use the above equations and solve them simultaneously as a system of equations.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (naudiapie):

Okay I see, noted